# SeNMo: A Self-Normalizing Deep Learning Model for Enhanced Multi-Omics Data Analysis in Oncology

### v1

@@ 2.1.1 Data Acquisition
The Cancer Genome Atlas (TCGA) houses one of the largest collections of high-dimensional multi-omics data sets for more than 33 different types of cancer for around 20,500 individual tumor samples [48]. The multi-omics data contains high-throughput RNA-Seq, DNA-Seq, miRNA-Seq, single-nucleotide variant (SNV ), copy number variation (CNV ), DNA methylation, and reverse phase protein array (RPPA) data [48]. Building cohorts for patient data lying across different data formats, modalities, and systems is not a trivial task. To curate the data and build the patient cohorts, we used our previously developed Multimodal Integration of Oncology Data System (MINDS) which is a metadata framework for fusing data from publicly available sources such as the TCGA’s Genomic Data Commons (GDC) and UCSC Xena portal [55] into a machine-learning ready format [28]. MINDS is available as open-source for the cancer research community and we have integrated MINDS into the SeNMo framework for enhanced outreach and benefit of researchers. We used pan-cancer data comprising 33 cancer types from TCGA and Xena for training, validation, and testing of our model. We used the CPTAC-LSCC data and Moffitt Cancer Center’s internal Lung SCC data for evaluating the model’s performance off-the-shelf. We also fine-tuned the model on the CPTAC-LSCC [56] and Moffitts’ LSCC [57] data to evaluate the model’s generalizability and transfer learning capabilities.
@@ 2.1.2 Data Modalities
Out of the 13 multi-omic modalities contained across each GDC-TCGA cancer database, the ones chosen were gene expression RNAseq, DNA methylation, miRNA stem-loop expression, reverse phase protein array (RPPA) data, DNA mutation, and clinical data. The rationale behind selecting these specific modalities was that these modalities have been frequently preferred over the other data types in studying cancer primarily due to their direct relevance in the fundamental processes of cancer progression, their diagnostic and prognostic capabilities, and their established technologies [58, 17]. These modalities directly influence and reflect key biological processes fundamental to cancer progression, making them extremely valuable for uncovering the molecular mechanisms driving the disease [58]. Moreover, the set of these modalities provides robust predictive and prognostic information, and their integration provides a holistic view of a tumor’s multi-omic profile [59, 58, 17]. Lastly, the selected modalities contained a static feature number across each cancer type that helped us in developing the standard data pre-processing pipeline for pan-cancer studies. Below we briefly describe each of the data modalities considered in this study, followed by the preprocessing steps undertaken to select the features used in training the SeNMo model.
@@
    DNA Methylation: DNA methylation is a key epigenetic modification where methyl groups are added to the DNA molecule, typically at cytosine bases adjacent to guanine, known as CpG sites [60]. This process plays a crucial role in regulating gene expression without altering the DNA sequence itself [60]. In the context of cancer, DNA methylation patterns are immensely valuable as a data modality. Aberrant methylation of certain genes can lead to their silencing or activation, contributing to oncogenesis and tumor progression [61]. By analyzing methylation profiles across different cancer types, researchers can identify diagnostic markers, predict disease progression, and tailor personalized treatment strategies [61]. Thus, DNA methylation serves as a critical biomarker in oncology, offering insights into the molecular mechanisms of cancer and enhancing the precision of therapeutic interventions [61]. DNA methylation is quantified through beta values, which range from 0 to 1, with higher values indicating increased methylation [62]. The beta values for DNA methylation data in TCGA-GDC were measured using the Illumina Human Methylation 450 platform, a sophisticated method for detailed methylation profiling [63]. The data consists of 485,576 unique cg and rs methylation sites across several tumor types [63].
@@
    Gene Expression RNAseq: Gene expression analysis via RNA sequencing (RNAseq) is a powerful data modality in cancer research, providing deep insights into the transcriptomic landscape of tumors [64]. This technique quantifies the presence and quantity of RNA in a biological sample at a given moment, allowing for a detailed view of transcriptional activity in a cell [64]. RNAseq helps to identify genes that are upregulated or downregulated in cancer cells compared to normal cells, offering clues about oncogenic pathways and potential therapeutic targets [65]. The data derived from RNAseq are crucial for understanding the molecular mechanisms driving cancer progression and for developing targeted treatments that can selectively affect cancerous cells without impacting normal tissue [65]. This modality not only supports the identification of biomarkers for diagnosis and prognosis but also aids in the personalization of cancer therapy, making it a cornerstone of precision oncology [65]. The TCGA-GDC gene expression data is derived from RNA sequencing (RNAseq), utilizing High-throughput sequence Fragments Per Kilobase of transcript per Million mapped reads (HTseq-FPKM) as a measure for normalization [66]. This approach normalizes raw reads counts by gene length and the number of mapped reads. Further processing includes incrementing the FPKM value by one and then applying a log transformation to stabilize variance and enhance statistical analysis [67]. The data spans 60,483 genes, with FPKM values indicating gene expression levels—values over 1000 signify high expression, whereas values between 0.5 and 10 indicate low expression [66, 68].
@@
    miRNA Stem Loop Expression: miRNA stem loop expression is a pivotal aspect in understanding the intricate regulatory mechanisms that miRNAs, or microRNAs, play in gene expression [69]. These small, non-coding RNA molecules typically function by binding to complementary sequences on target messenger RNA (mRNA) transcripts, leading to their silencing [69]. The expression of miRNAs involves a multi-step process, that ensures specific targeting and effective modulation of gene expression, crucial for both normal cellular function and pathological conditions, such as cancer, where miRNA expression profiles can markedly differ from those in healthy tissue [69]. The miRNA expression values for TCGA-GDC were measured using stem loop expression through illumina and values were transformed by adding one and being log transformed [70, 71]. These were mapped across 1880 features representing typical hsa-miRNA sites, where, like gene expression, expressions varied between high and low values.
@@
    Protein Expression: Reverse Phase Protein Array is an effective methodology that is similar to western blotting, and is used to quantify protein expression in tissue slides [72]. The technique involves transferring antibodies to a nitrocellulose-coated slide to bind specific proteins, forming quantifiable spots through a DAB calorimetric reaction and tyramide dye deposition, analyzed using "SuperCurve Fitting" software [72, 73]. This process allows for effective comparison of protein expression levels in tumor samples against benign samples, highlighting aberrant protein levels that drive the molecular phenotypes of cancer [72, 74]. Similarly, through the quantification of protein expression, RPPA uncovers the functional status of several signaling molecules, phosphorylation molecules, and metabolic molecules [73]. RPPA data has been generated from the profiling of nearly 500 antibody-proteins for a given patient and deposited in The Cancer Proteome Atlas (TCPA) portal [75]. Each data file contains 488 antigen ID (AGID), the peptide target ID, gene identifier that codes for the protein, various other identifiers, and the antigen’s corresponding expression level. Protein expression levels were normalized through log transformation and median centering after being calculated by the SuperCurve fitting software [76].
@@
    DNA Mutation:
    When a particular DNA is sequenced, oftentimes the first step is to determine which regions are mutated compared to a reference genome (also known as DNA-seq analysis) [77]. These differences in alignments are subsequently captured in a variant calling format (VCF) file, which encapsulates the characteristics (chromosome, position, base mutated, other features, etc.) of the differing regions [78]. When VCF files are further aggregated to remove low quality variants and only include somatic mutations, they produce Mutation Annotation Format (MAF) format files [79]. MAF files also only consider the most affected reference for comparison, while VCF files consider all reference transcripts [77]. Like VCF files, MAF files include characteristics of the mutations such as the start position, end position, reference allele, type of mutation (deletion, insertion, polymorphism), but also include many different quantifiable scores that assess the translational impact of the mutation with references to other databases that provide the clinical significance [79]. Such information is of high relevance since mutations of clinical significance usually contain large defects in protein structure, and this greatly affects the protein’s downstream function. This dysregulation ultimately makes such mutations the driving force behind the development of several cancers [80]. SIFT and PolyPhen are features within MAF files that quantify a mutation’s effect on its encoded protein and can be further analyzed. MAF files from TCGA-GDC each have a varying number of rows, as different samples have a different number of mutations, but each row contains 126 mutational characteristics [79]. Analyzing MAF files not only aids in the identification of biomarkers for diagnosis and prognosis but also supports the personalization of targeted cancer therapy, thereby establishing it as a cornerstone of precision oncology [58, 81].
@@
    Clinical Data: Clinical data plays a crucial role in cancer research, serving as the foundation for correlating biological data with patient outcomes and demographics [4]. The clinical data encompasses detailed patient information, which is instrumental in understanding the epidemiology of cancer, evaluating treatment responses, and improving prognostic assessments [4]. Integrating clinical data with genomic and proteomic analyses can uncover relationships between molecular profiles and clinical manifestations of cancer [9]. Among the many clinical features and phenotypes available, age, gender, race, and cancer stage are particularly emphasized in cancer research due to their significant impact on disease presentation, progression, and treatment efficacy [82, 83, 84, 85]. Age is a critical factor as the incidence and type of cancer often vary significantly with age, influencing both the biological behavior of tumors and the overall prognosis of patients [82]. Gender is another key determinant, as certain cancers are gender-specific, while others may show differences in occurrence and outcome between men and women, likely due to biological, hormonal, and social factors [83]. Race has been linked to differences in cancer susceptibility, mortality rates, and treatment outcomes, reflecting underlying genetic, environmental, and socioeconomic factors [84]. Finally, cancer stage at diagnosis is paramount for determining the extent of disease and guiding treatment decisions, directly correlating with survival rates [85]. These four features are prioritized because they can be consistently measured, provide fundamental insights into the patient’s condition, and are predictive of clinical outcomes, making them indispensable in both research and clinical settings.
@@ 2.1.3 Pre-processing
Multiomics data integrates diverse biological data modalities such as genomics, transcriptomics, proteomics, and metabolomics, to understand the complex mechanisms of diseases like cancer. However, before integration, this data requires multiple preprocessing steps to overcome the big p, small n problem and other associated challenges of high-throughput molecular data. Some of the most significant challenges include, (i) data heterogeneity, where each type of data encompasses unique properties and scales, (ii) volume and complexity, where overwhelming volume of data (often in terabytes), managing, storing, and processing requires substantial computational resources and advanced data management strategies, (iii) quality and variability incurred because of the different platforms resulting in batch effects, differing levels of sensitivity, noise, missingness, and varying error rates, and (iv) lack of standardization in how data is collected and processed across different laboratories and studies. These challenges are further pronounced when selecting the preprocessing steps to make the data machine learning-ready. Some of the tasks to consider while dealing with multi-omic data include:
@@
    Normalization and Scaling. Because of the diverse nature, each type of omics data requires specific normalization techniques to adjust for factors. (i.e. gene length in RNA-seq data or protein abundance in proteomics). Choosing the right normalization method based on the type of data is crucial to ensure that data is comparable [86, 87, 88].
@@
    Handling Missing Data. Multiomics datasets often contain missing values due to detection limits or experimental errors. At times, as in our case, an entire data modality for a patient is missing. Selecting the robust imputation method for missing data is critical to avoid biased interpretations. Typical imputation methods suggest using mean, median, KNN, Gaussian mixture clustering, Bayesian, and deep learning-based (autoencoders) techniques to handle imputations [89].
@@
    Dimensionality Reduction. The high dimensionality of multiomics data often exceeds the number of samples available, leading to the risk of overfitting. Techniques like principal component analysis (PCA), t-distributed stochastic neighbor embedding (t-SNE), features selection, features engineering and others are used to reduce dimensionality while preserving the most informative features of the data [90].
@@
    Data Annotation and Metadata. Proper annotation and comprehensive metadata are essential for the effective preprocessing of multiomics data. Metadata must capture details about sample collection, processing protocols, and experimental conditions, which are crucial for accurate data interpretation and reproducibility [91].
@@
    Integration Techniques. Integrating diverse datasets involves sophisticated statistical and computational methods. Techniques such as concatenation, transformation, and advanced modeling (ML/ DL algorithms) are usually used to merge these datasets coherently [92].
Addressing these challenges requires interdisciplinary expertise, including bioinformatics, statistics, and domain-specific knowledge. The pan-cancer multiomics data comes with intra- and inter-dataset correlations, heterogeneous measurement scales, missing values, technical variability, and other background noise. Here we describe the preprocessing steps that were used across molecular data modalities.
@@
  • Drop constant features. Next, constant/quasi constant features with a threshold of 0.998 were filtered out using Feature-engine, a A Python library for feature engineering and selection [93]. This eliminated features with no expression at all across every sample along with features that were noise, since the expression value was the same across every sample.
  • Remove duplicates features. Next, duplicate features between genes were identified that contained the same values across two seperate genes, and one of the genes were kept. This may reveal gene-gene relationships between the two genes stemming from a up-regulation pathway or could simply reflect noise.
  • Remove colinear features. Next, we filtered the features having low variance ($\approx$0.25) because we the features having high variance hold the maximum amount of information [94]. We used VarianceThreshold feature selector of scikit learn library that removes low-variance features based on the defined threshold [95]. We chose a threshold for each data modality so that the resulting features have the matching dimensions, as shown in Figure 2.
  • Remove low-expression genes. The gene expression data originally contained 60,483 features, with FPKM transformed numbers ranging from 0 to 12. Roughly 30,000 genes remained after the above-mentioned preprocessing steps, which was still a very high number of features. High expression values reveal important biological insights due to an indication that a certain gene product is transcribed in large quantities, revealing that gene features with large expression values within the dataset are highly relevant. Genes containing an expression value greater than 7 (127 FPKM value) were kept, while the rest were discarded. Around 3000 genes remained after this process, all of which ranged from values between 7 and 12.
  • Handle missing features. We handled missing features at two levels of data integration separately. For the features within each modality and within each cancer type, the missing values were imputed with the mean of the samples for that feature. This resulted in the full-length features vector for each sample. Across different cancers and across different modalities, we padded the missing features with zeros. One may opine that this is equivalent to zero-padding prevalent in the bio-statistics, but we argue that padding zeros across cancers and modalities is not imputation when integrating very high dimensional, and high-sample sized data. Moreover, zero-padding has also been shown to improve the quality of gene studies results across disease data including breast cancer [96]. In deep learning, zero imputation technique shows the best performance compared to other imputation techniques and deficient data removal techniques [97, 98]. Moreover, there is a line of work that simply used zero padding by to minimize the noise in data and achieved state-of-the-art performance on respective datasets [99, 100, 101]

### v2

@@ 2.1.1 Data Acquisition
TCGA houses one of the largest collections of high-dimensional multi-omics datasets, comprising over 20,500 individual tumor samples from 33 different cancer types [49]. The available data includes high-throughput RNA sequencing (RNA-Seq), DNA sequencing (DNA-Seq), microRNA sequencing (miRNA-Seq), single nucleotide variants, copy number variations, DNA methylation, and reverse-phase protein array (RPPA) data [49]. Building cohorts from this diverse data, spanning multiple formats, modalities, and systems, presents significant challenges. To curate and establish patient cohorts, we utilized our previously developed Multimodal Integration of Oncology Data System (MINDS), a metadata framework designed to fuse data from publicly available sources like TCGA-GDC and UCSC Xena Portal into a machine learning-ready format [49, 56, 29]. MINDS is freely accessible to the cancer research community, and has been integrated into the SeNMo framework to enhance its usability and benefit to researchers. For training, validation, and testing, we used pan-cancer data from TCGA and Xena, covering 33 cancer types, as summarized in Table 1. We further fine-tuned the model using data from the CPTAC-LSCC [57] and Moffitt’s LSCC datasets [58] to evaluate the generalizability and transfer learning capabilities of SeNMo.
@@ 2.1.2 Data Modalities
From the 13 available multi-omic modalities present in each cancer dataset, we selected gene expression (RNAseq), DNA methylation, miRNA stem-loop expression, RPPA data, DNA mutation, and clinical data. These modalities were chosen based on their frequent use in cancer studies due to their direct relevance to the fundamental processes of cancer progression, as well as their diagnostic and prognostic capabilities [59, 17]. They offer direct insights into key biological processes fundamental to cancer progression, making them extremely valuable for uncovering the molecular mechanisms driving the disease [59]. Furthermore, these selected modalities provide robust predictive and prognostic information, and their integration gives a holistic view of a tumor’s multi-omic profile [60, 59, 17]. Importantly, each modality had a consistent number of features across all cancer types, which facilitated the development of a standardized data preprocessing pipeline for pan-cancer studies. Below is a brief description of each data modality considered in this study, followed by the preprocessing steps used to select features for training the SeNMo model.
@@
    DNA methylation: DNA methylation is an epigenetic modification involving the addition of methyl groups to the DNA molecule, typically at cytosine bases adjacent to guanine, known as CpG sites [61]. This modification plays a crucial role in regulating gene expression without altering the DNA sequence [61]. In cancer, aberrant methylation can lead to the silencing or activation of genes, contributing to oncogenesis and tumor progression [62]. Analyzing methylation profiles across different cancer types helps identify risk and diagnostic markers, predict disease progression, and support personalized treatment strategies [62]. DNA methylation is quantified through beta values ranging from 0 to 1, with higher values indicating increased methylation [63]. The beta values for TCGA-GDC methylation data were obtained using the Illumina Human Methylation 450 platform, which provides detailed methylation profiling [64]. The dataset contains 485,576 unique cg and rs methylation sites across multiple tumor types [64].
@@
    Gene expression (RNAseq): Gene expression analysis through RNA sequencing (RNAseq) is a powerful modality in cancer research, providing insights into the transcriptomic landscape of tumors [65]. This technique quantifies the presence and quantity of RNA in a biological sample, giving a detailed view of transcriptional activity in a cell [65]. RNAseq helps identify genes that are upregulated or downregulated in cancer cells compared to normal cells, offering clues about oncogenic pathways and potential therapeutic targets [66]. TCGA-GDC gene expression data was obtained from RNAseq, utilizing High-throughput sequence Fragments Per Kilobase of transcript per Million mapped reads (HTseq-FPKM) for normalization [67]. This approach normalizes raw read counts by gene length and the number of mapped reads, with further processing involving incrementing the FPKM value by one followed by log transformation to stabilize variance and enhance statistical analysis [68]. The dataset includes 60,483 genes, with FPKM values indicating gene expression levels. Values above 1000 signify high expression, while values between 0.5 and 10 indicate low expression [67, 69].
@@
    miRNA stem loop expression: miRNA stem-loop expression plays a pivotal role in understanding the regulatory mechanisms of miRNAs (microRNAs) in gene expression [70]. miRNAs are small, non-coding RNA molecules that function by binding to complementary sequences on target mRNA transcripts, leading to silencing [70]. The expression of miRNAs involves multiple steps to ensure specific targeting and effective modulation of gene expression, which is crucial for normal cellular function as well as pathological conditions like cancer [70]. miRNA expression values for TCGA-GDC were measured using stem-loop expression through Illumina, and values were log-transformed after the addition of one [71, 72]. The data represents 1880 features across hsa-miRNA sites, with expression levels varying between high and low.
@@
    Protein expression: Reverse Phase Protein Array (RPPA) is a laboratory technique similar to western blotting, used to quantify protein expression in tissue samples [73]. The method involves transferring antibodies onto nitrocellulose-coated slides to bind specific proteins, forming quantifiable spots via a DAB calorimetric reaction and tyramide dye deposition, analyzed using "SuperCurve Fitting" software [73, 74]. RPPA effectively compares protein expression levels in tumor and benign samples, highlighting aberrant protein levels that define the molecular phenotypes of cancer [73, 75]. RPPA data in TCGA was derived from profiling nearly 500 antibody-proteins for each patient and deposited in The Cancer Proteome Atlas portal [76]. Each dataset includes the antigen ID, peptide target ID, gene identifier that codes for the protein, and antigen expression levels. Protein expression levels were normalized through log transformation and median centering after being calculated by SuperCurve fitting software [77].
@@
    DNA mutation: Analyzing DNA sequences involves identifying mutated regions compared to a reference genome, resulting in Variant Calling Format (VCF) files detailing these differences [78, 79]. Aggregating VCF files to exclude low-quality variants and include only somatic mutations produces Mutation Annotation Format (MAF) files [80]. Unlike VCF files, which consider all reference transcripts, MAF files focus on the most affected references and include detailed characteristics and quantifiable scores that assess a mutation’s translational impact and clinical significance [80]. This information is critical because clinically significant mutations often result in major defects in protein structure, severely impacting downstream functions and contributing to cancer development [81]. The MAF files from TCGA-GDC contain 18,090 mutational characteristics [80].
@@
    Clinical data: Clinical and patient-level data play a crucial role in cancer research, providing the foundation for identifying and characterizing patient cohorts [4]. Clinical data includes detailed patient information that is instrumental in understanding cancer epidemiology, evaluating treatment responses, and improving prognostic assessments [4]. Integrating clinical data with genomic and proteomic analyses can uncover relationships between molecular profiles and clinical manifestations of cancer [9]. Key clinical and patient-level covariates such as age, gender, race, and disease stage are particularly important in cancer research due to their impact on disease presentation, progression, and treatment efficacy [82, 83, 84, 85]. Age is a critical factor as cancer incidence and type often vary significantly with age, influencing both the biological behavior of tumors and patient prognosis [82]. Gender also plays an important role, with certain cancers being gender-specific and others differing in occurrence and outcomes between genders due to biological, hormonal, and social factors [83]. Race and ethnicity are linked to differences in cancer susceptibility, mortality rates, and treatment outcomes, which reflect underlying genetic, environmental, and socioeconomic factors [84]. Finally, cancer stage and histology at diagnosis are paramount for determining disease extent, guiding treatment decisions, and correlating directly with survival rates [85].
@@ 2.1.3 Pre-processing
Multiomics data integrates diverse biological data modalities such as genomics, transcriptomics, proteomics, and metabolomics, to understand the complex mechanisms of diseases like cancer. However, before integration, this data requires multiple preprocessing steps to overcome the big P, small n problem and other associated challenges of high-throughput molecular data. The big P, small n problem refers to a large number of features (P) and a small number of samples (n) in the data [86]. The pan-cancer multi-omics data comes with intra- and inter-dataset correlations, heterogeneous measurement scales, missing values, technical variability, and other background noise. Key challenges include: (i) data heterogeneity, where each data type has unique properties and scales, (ii) volume and complexity, which involve managing and processing overwhelming volumes of data, often in terabytes, (iii) quality and variability, which stem from different platforms causing batch effects, sensitivity differences, noise, varying error rates, and missingness, and (iv) lack of standardization in data collection and processing across laboratories and studies. These challenges complicate the preprocessing needed to make the data machine learning-ready. The key preprocessing tasks for multi-omic data are:
@@
    Normalization and scaling. Due to their diverse nature, each omics data type requires specific normalization techniques (e.g., gene length adjustment in RNA-seq or protein abundance correction in proteomics). Choosing the right normalization method ensures that data are comparable across modalities [87, 88, 89].
@@
    Handling missing data. Multiomics datasets often contain missing values due to detection limits or experimental errors. In some cases, an entire data modality for a patient may be missing. Robust imputation methods are critical to avoid biased interpretations. Common methods include mean, median, kNN, Gaussian mixture clustering, Bayesian approaches, and deep learning-based techniques such as autoencoders [90].
@@
    Dimensionality reduction. The high dimensionality of multi-omics data often exceeds the number of samples available, increasing the risk of overfitting. Techniques like principal component analysis, t-distributed stochastic neighbor embedding, feature selection, and feature engineering are used to reduce dimensionality while preserving the most informative aspects of the data [91].
@@
    Data annotation and metadata. Proper annotation and comprehensive metadata are essential for effective preprocessing of multiomics data. Metadata should capture details about sample collection, processing protocols, and experimental conditions to ensure accurate data interpretation and reproducibility [92].
@@
    Integration techniques. Integrating diverse datasets involves sophisticated statistical and computational methods. Techniques such as concatenation, transformation, and advanced modeling (e.g., machine learning or deep learning algorithms) are typically used to merge these datasets coherently [93].
Addressing these challenges requires interdisciplinary expertise, including bioinformatics, statistics, and domain-specific knowledge. Here, we describe the preprocessing steps used across molecular data modalities.
@@
  • Drop constant features. Next, constant/quasi-constant features with a threshold of 0.998 were filtered out using Feature-engine, a Python library for feature engineering and selection [94]. This eliminated features with no expression at all across every sample along with features that were noise, since the expression value was the same across every sample.
  • Remove duplicates features. Next, duplicate features between genes were identified that contained the same values across two seperate genes, and one of the genes was kept. This may reveal gene-gene relationships between the two genes stemming from an up-regulation pathway or could simply reflect noise.
  • Remove colinear features. Next, we filtered the features having low variance ($\approx$0.25) because the features having high variance hold the maximum amount of information [95]. We used VarianceThreshold feature selector of scikit learn library that removes low-variance features based on the defined threshold [96]. We chose a threshold for each data modality so that the resulting features have matching dimensions, as shown in Figure 2.
  • Remove low-expression genes. The gene expression data originally contained 60,483 features, with FPKM transformed numbers ranging from 0 to 12. Roughly 30,000 genes remained after the above-mentioned preprocessing steps, which was still a very high number of features. High expression values reveal important biological insights due to an indication that a certain gene product is transcribed in large quantities, revealing that gene features with large expression values within the dataset are highly relevant. Genes containing an expression value greater than 7 (127 FPKM value) were kept, while the rest were discarded. Around 3,000 genes remained after this process, all of which ranged from values between 7 and 12.
  • Handle missing features. We handled missing features at two levels of data integration. First, for the features within each modality and cancer type, the missing values were imputed with the mean of the samples for that feature. This resulted in the full-length feature vector for each sample. Second, across different cancers and modalities, we padded the missing features with zeros. One may opine that this is equivalent to zero-padding prevalent in the bio-statistics, but we argue that padding zeros across cancers and modalities is not an imputation when integrating very high dimensional, and high-sample-sized data. In deep learning, the zero imputation technique shows the best performance compared to other imputation techniques and deficient data removal techniques [97, 98]. Moreover, there is a line of work that simply used zero padding to minimize the noise in data and achieved state-of-the-art performance on respective datasets [99, 100].